OpenStudy (anonymous):

What is f(-3)? on relations and functions A:-2 B:-1 C:1 D:5

jimthompson5910 (jim_thompson5910):

Are you given a graph?

OpenStudy (anonymous):

No

jimthompson5910 (jim_thompson5910):

what other information about f(x) are you given?

jimthompson5910 (jim_thompson5910):

are you given a chart?

OpenStudy (anonymous):

Thats all that is in the question but i found the answer now i have a different problem.

OpenStudy (anonymous):

Given f(x) = x – 10, what is f(–7)?

jimthompson5910 (jim_thompson5910):

oh ok, f(-7) means that you replace every copy of x with -7 and evaluate

jimthompson5910 (jim_thompson5910):

f(x) = x – 10 f(-7) = -7 – 10 f(-7) = ??

OpenStudy (anonymous):

-17

jimthompson5910 (jim_thompson5910):

good

OpenStudy (anonymous):

thanks i have one more on these types of problems

OpenStudy (anonymous):

by the way im gonna be your fan

jimthompson5910 (jim_thompson5910):

ok go ahead and post

OpenStudy (anonymous):

Given f(x)=-x+15 what is f(-3)

jimthompson5910 (jim_thompson5910):

same thing, just with a different value

jimthompson5910 (jim_thompson5910):

replace every copy of x with -3 and evaluate

OpenStudy (anonymous):

-18

jimthompson5910 (jim_thompson5910):

you replace x, not -x

jimthompson5910 (jim_thompson5910):

f(x) = -x + 15 f(x) = -1*x + 15 f(-3) = -1*(-3) + 15 f(-3) = ???

OpenStudy (anonymous):

18

jimthompson5910 (jim_thompson5910):

better

OpenStudy (anonymous):

thanks i do online school so my grade was an 80% witch is passing